How do you solve for x in the following problem: 5+(1/x-2)=35?
I've tried distributing all the way where I'm left with 1/x= 32 but I am getting confused on how to get the x by itself.

Answer by Cromlix(4381)

5 + 1/x - 2 = 35
Common denominator = x - 2
5(x - 2) + 1 = 35(x - 2)
5x - 10 + 1 = 35x - 70
Collect like terms:
-30x = - 61
30x = 61
x = 61/30
x = 2 1/30
Hope this helps.
:-)